During the manufacture of cement for the construction industry limestone, CaCO3

(s), is decomposed to quicklime, CaO, by vigorous heating. What is the percentage yield and percent difference of the reaction when 13.1 tonne (1 Mg = 1 tonne=106g) of quicklime are obtained from the treatment of 24.5 tonne of limestone?

CaCO3(s) → CaO(s) + CO2(g)


1
Expert's answer
2020-10-01T07:52:42-0400
"n(CaCO_3) = m/M ={24.5*10^6g \over 100g/mol} = 24.5*10^4 mol"

"n(CaCO_3) = n(CaO) = 24.5*10^4mol = 2.45*10^5mol"

From data:


"n(CaO) = {13.1*10^6g \over 56g/mol} = 2.34*10^5mol"

Then:


"\eta = {2.34*10^5 \over 2.45*10^5} = 0.955 = 95.5\%"